OpenStudy (jennychan12):

Find the limit. Use L'Hopital's rule if possible. Use a more elementary method if possible. If L'Hopital doesn't apply, explain why. lim x-->0 (x^(x^2))

OpenStudy (jennychan12):

\[\lim_{x \rightarrow 0} x^{x^2}\]

OpenStudy (jennychan12):

@agent0smith can you help?

OpenStudy (anonymous):

don't know how to do with without l'hopital

OpenStudy (jennychan12):

it is a l'hopital question

OpenStudy (tkhunny):

Have you considered a logarithm?

OpenStudy (jennychan12):

yes, i took ln of both sides

OpenStudy (anonymous):

take the log, take the limit of the log, then exponentiate

OpenStudy (tkhunny):

?? There's only one side.

OpenStudy (anonymous):

or rewrite as \[x^{x^2}=e^{x^2\ln(x)}\] then take the limit of that one the work is all the same, finding \[\lim_{x\to 0}x^2\ln(x)\] which you usually do by rewriting as \[\lim_{x\to 0}\frac{\ln(x)}{\frac{1}{x^2}}\] to get it in the form \(\frac{\infty}{\infty}\) so you can use l'hopital

OpenStudy (jennychan12):

ok, thanks :)

OpenStudy (agent0smith):

You can somewhat evaluate it analytically, since the x^2 will approach 0 faster than the x, so it'll approach 1. Wolfram uses a similar method to @satellite73 http://www4a.wolframalpha.com/Calculate/MSP/MSP4614217a9acbb15hadbg0000230hc10ec765i6f0?MSPStoreType=image/png&s=31&w=492&h=938